# Fixed-relative-switch strategies for learning based event-triggered control of nonlinear multiagent systems

### v2

@@ III The Controller Design
$Step\text{ }n$: Differentiating ${{z}_{i,n}}$, one gets
${{\dot{z}}_{i,n}}=$ $u_{i}+\hat{W}_{i,n}^{T}{{E}_{i,n}}\left({{{\hat{\bar{x}}}}_{i,n}}\right)+{{q}_{i,n}}{{\psi}_{i,1}}+{{\hat{\varpi}}_{i,n}}(t)-{{\dot{\bar{\alpha}}}_{i,n}}$ (34)
Choose the Lyapunov function as
@@
where ${{\eta}_{i,n}}$ denotes a designed positive parameter.
By utilizing RBF NNs, one has
${{\bar{Z}}_{i,n}}\left({{T}_{i,n}}\right)=$ $K_{i,n}^{*T}{{E}_{i,n}}\left({{T}_{i,n}}\right)+{{\delta}_{i,n}}\left({{T}_{i,n}}\right)$
$\left|{{\delta}_{i,n}}\left({{T}_{i,n}}\right)\right|\leq$ $\bar{\delta}_{i,n}$ (36)
where $\bar{\delta}_{i,n}$ denotes an unknown positive parameter.
The control law and adaptive parameters are constructed as
@@
where ${{r}_{i,n}}$, ${{h}_{i,n}}$ and ${{\lambda}}_{i}$ are designed parameters.
@@ III-A Fixed-threshold strategy
In this strategy, the adaptive event-triggered controller is reformulated as
@@
The triggering condition is defined as
@@
where $\vartheta_{i}(t)=w_{i}(t)-u_{i}(t)$ represents the measurement-triggered error, while $t_{s}$ indicates the controller’s update time, with $s\in Z$. The parameters $\bar{\pi}_{i}$, $\pi_{i}$ and $\mu_{i}$ are designed and positive, $\bar{\pi}>\pi_{i}$. The control signal ${{u}_{i}}\left({{t}_{s+1}}\right)$ is applied to the system when condition (42) is activated. During the interval $\text{t}\in\left[{{t}_{s}},{{t}_{s+1}}\right)$, i.e. $\left|{{w}_{i}}\left(t\right)-{{u}_{i}}\left(t\right)\right|<{{\pi}_{i}}$, the controller maintains a constant value of ${{w}_{i}}\left({{t}_{s}}\right)$. For a function ${{\epsilon}_{i}}{{\left(t\right)}}$ that changes continuously over time that satisfies ${{\epsilon}_{i}}\left({{t}_{s}}\right)=0$ and ${{\epsilon}_{i}}\left({{t}_{s+1}}\right)=\pm 1$ with $\left|{{\epsilon}_{i}}\left(t\right)\right|\leq 1$, it follows that ${{w}_{i}}\left(t\right)={{u}_{i}}\left(t\right)+{{\epsilon}_{i}}\left(t\right){{\pi}_{i}}$.
Based on Lemma 2, the inequality $-{{\epsilon}_{i}}(t){{\pi}_{i}}{{z}_{i,n}}-{{\bar{\pi}}_{i}}{{z}_{i,n}}\tanh(\frac{{{z}_{i,n}}{{{\bar{\pi}}}_{i}}}{{{\mu}_{i}}})\leq 0.2875{{\mu}_{i}}$ holds, then, differentiate ${{V}_{i,n}}$ by utilizing the Young’s inequality and (36)-(38), one gets
@@
where ${{\iota}_{i,n}}=\frac{c_{i,n}^{2}}{2}+\frac{\bar{\delta}_{i,n}^{2}}{2}+\frac{\kappa_{i,n}^{2}\sigma_{0,n}^{2}}{2}+0.2785{{\mu}_{i}}$, $c_{i,n}$ is a positive parameter.
@@ III-B Relative-threshold strategy
In the fixed-threshold strategy, it is observed that the threshold $\pi_{i}$ remains constant regardless of the control signal’s magnitude. Nevertheless, when addressing a stabilization problem, adopting a variable threshold for the triggering condition becomes advantageous, according to the approach in [21][22]. Specifically, when the control signal $u_{i}$ is huge, a greater error of measurement can be tolerated, allowing for longer update intervals. Conversely, as the system states stabilize towards equilibrium, with $u_{i}$ approaching zero, a smaller threshold enables more accurate control, enhancing overall system performance. Then, it is proposed the following adaptive event-triggered controller:
@@
where $t_{s}$ represents the time when the controller is updated, $s\in Z$, $\mu_{i}$, $\Delta_{i}$, $10<\Delta_{i}<$ 1, $\pi^{*}_{i}>0$, and $\bar{\pi}^{*}_{i}>\pi^{*}_{i}/(1-\Delta_{i})$ are all positive designed parameters.
From (49), we have $w_{i}(t)=(1+\rho_{1,i}(t))u_{i}(t)+\rho_{2,i}(t)\pi_{i}^{*}$ in the interval $[t_{s},t_{s+1}]$, where $\rho_{1,i}(t)$ and $\rho_{2,i}(t)$ denote time-varying parameters satisfying $|\rho_{1,i}(t)|\leq 1$ and $|\rho_{2,i}(t)|\leq 1$.
Thus, one has
@@
Since $\forall a_{i}\in R$ and $\mu_{i}>0$, from (45) we get $z_{i,n}w_{i}(t)\leq 0$, one has
$\frac{z_{i,n}w_{i}(t)}{1+\rho_{1,i}(t)\Delta_{i}}\leq\frac{z_{i,n}w_{i}(t)}{1+\Delta_{i}}$ (48)
$\frac{\rho_{2,i}\pi_{i}^{*}}{1+\rho_{1,i}(t)\Delta_{i}}\leq\frac{\pi_{i}^{*}}{1-\Delta_{i}}$ (49)
@@
III-C Switch-threshold strategy
In this portion, it is introduced a switch-threshold strategy. The relative-threshold approach adjusts the threshold based on the control signal’s magnitude, allowing for longer update intervals when the signal is huge and more precise control as the signal nears zero, thus improving performance. However, a very large control signal can lead to significant measurement errors and abrupt changes during updates. In contrast, the fixed-threshold strategy maintains a consistent upper limit on measurement errors, regardless of signal size. Considering these aspects, we propose the following switched-threshold strategy. The triggering condition is defined as follows:
@@ III-C Switch-threshold strategy
where $G$ is the designed switching gate.
For this switch-threshold strategy, one has
@@
Since the switch-threshold strategy employs an identical control law to both the fixed-threshold and relative-threshold strategies, the ultimate bounds for tracking and stabilization errors remain consistent with those of the two previously mentioned strategies.
IV Stability Analysis
Considering the previous analysis, we have developed observers and designed an adaptive consensus tracking controller via RBF NNs, backstepping method and three event-triggered strategies. In this section, we demonstrate the boundedness of all signals by employing Lyapunov stability theory. In addition, the event-triggered strategies are proven to be effective, and the Zeno behavior is avoided.
@@ IV Stability Analysis
$Theorem\text{ 1}$: For the nonlinear MASs in equation (1) with the event-triggered controllers in equation (40)(44), if the initial condition satisfies $V(0)\leq\Omega$, the error signals $z_{i,k}$, $\tilde{c}_{i,k}$, ${{\tilde{W}}_{i,k}}$, ${{\tilde{\Theta}}_{i}}$, ${{e}_{i,k}}$, and ${{\psi}_{i,k}}$ are all uniformly bounded. Furthermore, the errors of consensus tracking between the followers’ outputs and the leader’s trajectory signal can be minimized to a specified range.
$Proof$: See Appendix A.
V Illustrative Example
@@
This section will examine an illustrative example to validate the performance and effectiveness of the proposed approach. The system comprises one leader agent and four follower agents. The dynamics of each agent are described by:
@@
The communication topology is illustrated in Fig. 1. The connection matrix linking the leader to the followers is represented as $\mathcal{B}=diag\{0,1,0,0\}$. And the adjacency matrix $\mathcal{A}$ and the Laplacian matrix $\mathcal{L}$ are presented as follows:
$\mathcal{A}=\left[\begin{array}[]{cccc}0&1&0&0\\ 0&0&0&0\\ 0&1&0&0\\ 1&0&0&0\\ \end{array}\right]$, $\mathcal{L}=\left[\begin{array}[]{cccc}1&-1&0&0\\ 0&0&0&0\\ 0&-1&1&0\\ -1&0&0&1\\ \end{array}\right]$.
The initial conditions of the four followers and state observers are selected as follows:
@@
Based on the parameter selection guidelines outlined in the stability analysis, the quantitative values of the design parameters are as follows: for fix-threshold strategy, ${{\pi}_{i}}=2.5$, ${{\bar{\pi}}_{i}}=4$, ${{\mu}_{i}}=5.4$, for relative-threshold strategy, ${{\pi}_{i}^{*}}=2$, ${{\bar{\pi}}_{i}^{*}}=4$, $\Delta_{i}=0.245$, for switch-threshold strategy $G=6$, for the first order low pass filter, $m_{i}=0.005$, for other parameters, ${{h}_{i,1}}={{h}_{i,2}}=50$, ${{r}_{i,1}}={{r}_{i,2}}=-100$, ${{c}_{i,1}}={{c}_{i,2}}=100$, ${{\eta}_{i,1}}={{\eta}_{i,2}}=0.01$, ${{q}_{i,1}}=350$, ${{q}_{i,2}}=0.5$, ${{\lambda}_{i}}=120$ and ${{o}_{i}}=25$.
In Fig. 2, shortly after the process begins, all followers within the system consistently follow the leader, successfully achieving the adaptive consensus control objective. Fig. 3 is presented to illustrate that the tracking error has converged to zero. The total triggering number for sample-data strategy is 5,000 times, and the triggering numbers for the three threshold strategies are detailed in Table I. Figs. 4 shows that the interval of events respectively triggered by three strategies. According to the statistics, it is clear that under the fixed-threshold strategy, the number of triggers is the highest. Under the relative-threshold strategy, the number of triggers is the lowest, while under the switch-threshold strategy, the number of triggers falls and both types of triggers play an indispensable role in this strategy.
@@
Figure 2: Consensus tracking performance.
Figure 3: Trajectories of tracking errors $z_{i,1}$($i=1,2,3,4.$)
@@
Switch-threshold strategy

### v3

@@ III Event-Triggered Controller Design
$Step\text{ }n$: Choose the Lyapunov function as
@@
where ${{\eta}_{i,n}}$ denotes a designed positive parameter. By RBF NNs, one has ${{\bar{Z}}_{i,n}}\left({{T}_{i,n}}\right)=K_{i,n}^{*T}{{E}_{i,n}}\left({{T}_{i,n}}\right)+{{\delta}_{i,n}}\left({{T}_{i,n}}\right)$, where $\bar{\delta}_{i,n}$ denotes an unknown positive parameter, with $\left|{{\delta}_{i,n}}\left({{T}_{i,n}}\right)\right|\leq\bar{\delta}_{i,n}$. With designed parameters ${{r}_{i,n}}$, ${{h}_{i,n}}$, ${{\lambda}}_{i}$, the control law and adaptive parameters are designed as
@@
We now develop three event-triggered control strategies that provably ensure consensus tracking for the MASs. Theoretical analysis is also provided for their convergence.
@@ III-A Fixed-threshold strategy
The adaptive event-triggered controller is reformulated as
@@
where the triggering condition is defined as
@@
where $\vartheta_{i}(t)=w_{i}(t)-u_{i}(t)$ is the measurement-triggered error and $t_{s}$ is the controller’s update time with $s\in Z$. The parameters $\bar{\pi}_{i}$, $\pi_{i}$ and $\mu_{i}$ are designed to be positive with $\bar{\pi}>\pi_{i}$. The control signal ${{u}_{i}}\left({{t}_{s+1}}\right)$ is applied to the system when condition (30) is activated. During the interval $\text{t}\in\left[{{t}_{s}},{{t}_{s+1}}\right)$, i.e. $\left|{{w}_{i}}\left(t\right)-{{u}_{i}}\left(t\right)\right|<{{\pi}_{i}}$, the controller maintains a constant value of ${{w}_{i}}\left({{t}_{s}}\right)$. For a function ${{\epsilon}_{i}}{{\left(t\right)}}$ that continuously changes over time and satisfies ${{\epsilon}_{i}}\left({{t}_{s}}\right)=0$ and ${{\epsilon}_{i}}\left({{t}_{s+1}}\right)=\pm 1$ with $\left|{{\epsilon}_{i}}\left(t\right)\right|\leq 1$, it follows that ${{w}_{i}}\left(t\right)={{u}_{i}}\left(t\right)+{{\epsilon}_{i}}\left(t\right){{\pi}_{i}}$. By Lemma 2, $-{{\epsilon}_{i}}(t){{\pi}_{i}}{{z}_{i,n}}-{{\bar{\pi}}_{i}}{{z}_{i,n}}\tanh(\frac{{{z}_{i,n}}{{{\bar{\pi}}}_{i}}}{{{\mu}_{i}}})\leq 0.2875{{\mu}_{i}}$ holds, then we differentiate ${{V}_{i,n}}$ by the Young’s inequality and (25)-(27), with ${{\iota}_{i,n}}=\frac{c_{i,n}^{2}}{2}+\frac{\bar{\delta}_{i,n}^{2}}{2}+\frac{\kappa_{i,n}^{2}\sigma_{0,n}^{2}}{2}+0.2785{{\mu}_{i}}$ and a positive parameter $c_{i,n}$, and get the result below:
@@ III-B Relative-threshold strategy
In the fixed-threshold strategy, the threshold $\pi_{i}$ remains constant regardless of the magnitude of the control signal. However, it is preferable to set a variable threshold for the triggering condition to achieve system stabilization [24]. Specifically, when the control signal $u_{i}$ is huge, a greater measurement error is tolerated, allowing for longer update intervals. Conversely, when $u_{i}$ approaches 00, the system states stabilize towards equilibrium and a smaller threshold enables more accurate control, improving the overall performance. Then, we propose the adaptive event-triggered controller:
@@
where $t_{s}$ represents the time when the controller is updated, $s\in Z$, $\mu_{i}$, $\Delta_{i}$, $10<\Delta_{i}<$ 1, $\pi^{*}_{i}>0$, and $\bar{\pi}^{*}_{i}>\pi^{*}_{i}/(1-\Delta_{i})$ are all positive designed parameters. From (34), we have $w_{i}(t)=(1+\rho_{1,i}(t))u_{i}(t)+\rho_{2,i}(t)\pi_{i}^{*}$ in the interval $[t_{s},t_{s+1}]$, where $\rho_{1,i}(t)$ and $\rho_{2,i}(t)$ denote time-varying parameters satisfying $|\rho_{1,i}(t)|\leq 1$ and $|\rho_{2,i}(t)|\leq 1$. Thus, one has
@@
III-C Switched-threshold strategy
We now introduce a switched-threshold strategy. The relative-threshold approach adjusts the threshold based on the control signal’s magnitude, allowing for longer update intervals when the signal is huge and more precise control as the signal nears zero, thus improving performance. However, a very large control signal can lead to significant measurement errors and abrupt changes during updates. In contrast, the fixed-threshold strategy maintains a consistent upper limit on measurement errors, regardless of signal size. Then we propose the control strategy whose switching gate $G$ and the triggering condition are defined below:
@@ III-C Switched-threshold strategy
With $t\in[t_{s},t_{s+1})$, we have
@@
Since the switched-threshold strategy employs an identical control law for both the fixed-threshold and relative-threshold strategies, the ultimate bounds for tracking and stabilization errors remain consistent with those of the above two.
In summary, we have established an adaptive consensus tracking control framework which integrates state-disturbance observer design, learning based dynamics approximation, backstepping method and three event-triggered control strategies. Finally, we prove that all error signals are bounded by employing Lyapunov stability theory. In addition, we also confirm the soundness of the proposed control strategies that effectively avoid Zeno behaviors.
Theorem 1.
@@
For the nonlinear MASs in equation (II-A) under the event-triggered controllers in equation (28)(32), if the initial condition of the MASs satisfies $V(0)\leq\Omega$, the error signals $z_{i,k}$, $\tilde{c}_{i,k}$, ${{\tilde{W}}_{i,k}}$, ${{\tilde{\Theta}}_{i}}$, ${{e}_{i,k}}$, and ${{\psi}_{i,k}}$ are all uniformly bounded. Furthermore, the errors of consensus tracking between the followers’ outputs and the leader’s trajectory signal can be minimized to a specified range.
Proof.
For the follower agents $i=1,...,N$, define the overall Lyapunov function of the MASs (II-A) as
Define the parameters $r_{i,1}$, $r_{i,2}$, $r_{i,k}$ and $r_{i,n}$ as ${{r}_{i,1}}=-\left({{d}_{i}}+{{b}_{i}}\right)+r_{i,1}^{*}$, ${{r}_{i,2}}=-\frac{{{d}_{i}}+{{b}_{i}}}{2}-1+r_{i,2}^{*}$, ${{r}_{i,k}}=-\frac{3}{2}+r_{i,k}^{*}$ and ${{r}_{i,n}}=-\frac{1}{2}+r_{i,n}^{*}$ with $k=3,...,n-1$. Based on the Young’s inequality, while $\left\|{{{\dot{\varpi}}}_{i,k}}\right\|\leq\kappa_{i,k}^{*}\left\|{{{\tilde{\tau}}}_{i,k}}\right\|$, one has
where $\kappa_{i,k}^{*}=(3-2{{\kappa}_{i,k}})/2$, $r_{i,k}^{*}$ and $\kappa_{i,k}^{*}$ are unknown parameters for stability analysis, with $k=1,...,n$. And $\hbar(\bullet)$ is the eigenvalue of the given matrix. Select the matrix ${{H}_{i}}$ such that ${{\hbar}_{\min}}[{{H}_{i}}-(1+\sum\limits_{k=1}^{n}{(\kappa_{i,k}^{2}+\kappa_{i,k}^{4})}){{I}_{n}}]={{\wp}_{i}}/2{{\hbar}_{\max}}({{F}_{i}})$, where ${{\wp}_{i}}$ is a positive parameter. Differentiate ${{e}_{i,2}}$ and ${{e}_{i,k+1}}$ with respect to time, one has ${{\dot{e}}_{i,2}}=-\frac{{{e}_{i,2}}}{{{m}_{i,2}}}+{\Gamma_{i,2}}$, ${{\dot{e}}_{i,k+1}}=-\frac{{{e}_{i,k+1}}}{{{m}_{i,k+1}}}+{\Gamma_{i,k+1}}$, where ${{\Gamma}_{i,2}}=-\frac{{{r}_{i,1}}{{{\dot{z}}}_{i,1}}}{{{d}_{i}}+{{b}_{i}}}+\frac{{{{\hat{\Theta}}}_{i}}}{2c_{i,1}^{2}\left({{d}_{i}}+{{b}_{i}}\right)}{{\dot{z}}_{i,1}}E_{i,1}^{T}\left({{T}_{i,1}}\right){{E}_{i,1}}\left({{T}_{i,1}}\right)+\frac{{{{\dot{z}}}_{i,1}}}{2\left({{d}_{i}}+{{b}_{i}}\right)}+\frac{{{{\hat{\Theta}}}_{i}}}{c_{i,1}^{2}\left({{d}_{i}}+{{b}_{i}}\right)}{{z}_{i,1}}E_{i,1}^{T}\left({{T}_{i,1}}\right){{\dot{E}}_{i,1}}\left({{T}_{i,1}}\right)$ and ${{\Gamma}_{i,k+1}}=\frac{{{\dot{e}}_{i,k}}}{{{m}_{i,k}}}+{{q}_{i,k}}{{\dot{\psi}}_{i,1}}+\frac{{{{\hat{\Theta}}}_{i}}}{2c_{i,k}^{2}}{{\dot{z}}_{i,k}}E_{i,k}^{T}\left({{T}_{i,k}}\right){{E}_{i,k}}\left({{T}_{i,k}}\right)-{{r}_{i,k}}{{\dot{z}}_{i,k}}+\frac{{{{\dot{z}}}_{i,k}}}{2}+\frac{{{{\hat{\Theta}}}_{i}}}{c_{i,k}^{2}}{{z}_{i,k}}E_{i,k}^{T}\left({{T}_{i,k}}\right){{\dot{E}}_{i,k}}\left({{T}_{i,k}}\right)$. Inspired by [23], we have
${{\mathbb{M}}_{i,k}}$ =$=\Bigg{\{}\sum\limits_{i=1}^{N}{\sum\limits_{k=1}^{n}{(z_{i,k}^{2}+\tilde{\tau}_{i,k}^{2}+\frac{1}{{{\eta}_{i,k}}}\tilde{W}_{i,k}^{T}{{{\tilde{W}}}_{i,k}})}}+\sum\limits_{i=1}^{N}{\frac{1}{{{o}_{i}}}\tilde{\Theta}_{i}^{2}}$
where ${{\mathbb{M}}_{i,k}}$ is compact in ${{R}^{\dim\left({{\mathbb{C}}_{i,k}}\right)}}$, there exists an inequality $\left\|{{\Gamma}_{i,k+1}}\right\|\leq{{L}_{i,k+1}}$, where $L_{i,k+1}$ denotes a positive parameter. Applying the Young’s inequality, one can get ${{\Gamma}_{i,k+1}}{{e}_{i,k+1}}\leq\frac{\Xi}{2}+\frac{L_{i,k+1}^{2}e_{i,k+1}^{2}}{2\Xi}$, where $\Xi$ is a positive parameter, with $k=1,...,n-1$. Select the parameters as $\frac{1}{{{m}_{i,2}}}=\frac{{{d}_{i}}+{{b}_{i}}}{2}+\frac{L_{i,2}^{2}}{2\Xi}+m_{i,2}^{*}$ and $\frac{1}{{{m}_{i,k+1}}}=\frac{1}{2}+\frac{L_{i,k+1}^{2}}{2\Xi}+m_{i,k+1}^{*}$, where $m_{i,k+1}^{*}$ denotes an unknown positive parameter, with $k=2,...,n-1$. With $l=2,...,n$ and $k=1,...,n$, define
$\gamma=$ $*\displaystyle\sum\limits_{i=1}^{N}{\sum\limits_{k=1}^{n}{({{\iota}_{i,k}+{\frac{{{h}_{i,k}}}{2{{\eta}_{i,k}}}W_{i,k}^{\text{*}T}W_{i,k}^{\text{*}}})}}}+$
$\sum\limits_{i=1}^{N}({\sum\limits_{k=1}^{n-1}{\frac{\Xi}{2}+{\frac{{{\lambda}_{i}}}{2{{o}_{i}}}{{\left(\Theta_{i}^{*}\right)}^{2}}}}})$ (43)
To ensure the stability of the entire MASs, it is crucial to appropriately select parameter values while adhering to the following conditions: $-r_{i,k}^{*}>0,\text{ }-\kappa_{i,k}^{*}>0,\text{ }{{h}_{i,k}}>0,\text{ }{{\wp}_{i}}>0,\text{ }{{\lambda}_{i}}>0$ and ${{m}_{i,l}^{*}}>0$. Then, we have
$\dot{V}\leq-\beta V+\gamma$ (44)
Consequently, as time progresses, all consensus tracking errors will converge to a compact set defined as =$\Im=\{{{\Upsilon}_{1}}|\left\|{{\Upsilon}_{1}}\right\|\leq\sqrt{2\gamma/\beta}\}$. This means that the tracking errors can be modified and reduced to an arbitrarily small range by increasing the parameter $\beta$. From (30), one has
${{\dot{\alpha}}_{i}}\left(t\right)={{\dot{w}}_{i,n+1}}-\frac{\bar{\pi}_{i}{{{\dot{z}}}_{i,n}}}{{{\cosh}^{2}}\left(\frac{{{z}_{i,n}}{{{\bar{\pi}}}_{i}}}{{{\mu}_{i}}}\right)}$ (46)
According to [26, 27, 28], one gets
$\frac{d}{dt}\left|{{\vartheta}_{i}}\left(t\right)\right|=sign\left({{\vartheta}_{i}}\left(t\right)\right){{\dot{\vartheta}}_{i}}\left(t\right)\leq\left|{{{\dot{w}}}_{i}}\left(t\right)\right|$ (47)
Based on the stability analysis, it is imperative that there exists a positive parameter $\varPi$ such that $|{{{\dot{w}}}_{i}}(t)|\leq\varPi$. Based on (29)(30), we have ${{\vartheta}_{i}}({{t}_{s}})=0$ and $\underset{{}}{\mathop{{{\lim}_{t\to{{t}_{s+1}}}}}}\,{{\vartheta}_{i}}(t)={{\pi}_{i}}$. Additionally, for the time interval $t\in[{{t}_{s}},{{t}_{s+1}})$, the lower bound for the inter-execution time is given by ${{t}^{*}}\geq{{\pi}_{i}}/{\varPi}$. Following the same analysis in the proof of fixed-threshold strategy, according to (33)(34), the relative-threshold and switch-threshold strategy satisfy ${{t}^{*}}\geq({\Delta_{i}|u_{i}(t)|+{\pi}_{i}^{*}})/{\varPi}$ and ${{t}^{*}}\geq{max\{{\pi}_{i},{\pi}_{i}^{*}\}}/{\varPi}$, respectively. Consequently, the Zeno behavior is proficiently avoided. ∎
IV Illustrative Example
@@
This section presents an illustrative example to validate the soundness and performance of our theoretical approach. The MASs under investigation comprise one leader and four followers, where the dynamics of each agent is described by:
@@
The communication topology is shown in Fig. 1. The connection matrix linking the leader to the followers is represented as $\mathcal{B}=diag(0,1,0,0)$. And the adjacency matrix $\mathcal{A}$ and the Laplacian matrix $\mathcal{L}$ are given as follows:
$\mathcal{A}=\left[\begin{array}[]{cccc}0&1&0&0\\ 0&0&0&0\\ 0&1&0&0\\ 1&0&0&0\\ \end{array}\right],\leavevmode\nobreak\ \mathcal{L}=\left[\begin{array}[]{cccc}1&-1&0&0\\ 0&0&0&0\\ 0&-1&1&0\\ -1&0&0&1\\ \end{array}\right]$
The initial conditions of both the four followers along with their state observers are selected as follows:
@@
Following the parameter selection guidelines outlined in the stability analysis, the values of the design parameters are set as: for fix-threshold strategy, ${{\pi}_{i}}=2.5$, ${{\bar{\pi}}_{i}}=4$, ${{\mu}_{i}}=5.4$, for relative-threshold strategy, ${{\pi}_{i}^{*}}=2$, ${{\bar{\pi}}_{i}^{*}}=4$, $\Delta_{i}=0.245$, for switched-threshold strategy $G=6$, for the first order low pass filter, $m_{i}=0.005$, for other parameters, ${{h}_{i,1}}={{h}_{i,2}}=50$, ${{r}_{i,1}}={{r}_{i,2}}=-100$, ${{c}_{i,1}}={{c}_{i,2}}=100$, ${{\eta}_{i,1}}={{\eta}_{i,2}}=0.01$, ${{q}_{i,1}}=350$, ${{q}_{i,2}}=0.5$, ${{\lambda}_{i}}=120$ and ${{o}_{i}}=25$.
Figure 2: Consensus tracking performance.
Figure 3: Trajectories of tracking errors $z_{i,1}$($i=1,2,3,4.$)
In Fig. 2, shortly after the process begins, all followers within the system consistently follow the leader, successfully achieving adaptive consensus tracking. Fig. 3 illustrates that the tracking error has converged to zero. The total triggering number for the sample-data strategy is 5,000 times, and the triggering numbers for the three threshold strategies are detailed in Table I. Figs. 4 shows the interval of events respectively triggered by the three strategies. The statistics demonstrates that the highest number of triggers arises from the fixed-threshold strategy. Under the relative-threshold strategy, the number of triggers is the lowest, while under the switched-threshold strategy, the number of triggers falls and both types play an indispensable role in this strategy.
@@
Switched-threshold strategy